what is the equation of the line that passes through the points (-2,2)and (0,5)?
Answer by Lal(16)

(-2,2) and (0,5)

Find the slope, and then use the point-slope form to put it into 'y = mx + b' form.

For slope: change of y over change in x.
m = 5 - 2 / 0 + 2
m = 3/2

Use point-slope form with one of your points.
y - y1 = m( x - x1)
y - 2 = 3/2( x + 2)
y - 2 = 3/2x + 3
y = 3/2x + 5

The equation is: y = 3/2x + 5